OpenStudy (anonymous):

how would i write out this problem for which of these equations is y a function of x a. y=3x to the 3rd power b.xy=1 c.yto 2 nd power d. y to 3rd power=8x

OpenStudy (anonymous):

I not sure you write, I think a) y= (3x)^3

OpenStudy (anonymous):

yes thats how its written but is y a function of x how do i determine that

OpenStudy (anonymous):

you want find x and y intercept?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok: you plug x=0 find y plug y=0 find x

OpenStudy (anonymous):

Example: find y plug x=0 a) y= (3x)^3 y= (3*0)^3 y= 0 find x: plug y= 0 0 = (3x)^3 0 = 27(x)^3 0/27=x^3 x=0 (0,0)

OpenStudy (anonymous):

so is this equation y a function of x